# IBD-PSC: Input-level Backdoor Detection via Parameter-oriented Scaling Consistency

### v2

@@ 5.2 Main Results
As shown in Table 1-3, IBD-PSC consistently achieves promising performance in all cases across various datasets. For instance, it achieves AUROC and F1 scores approaching 1.0, indicating its effectiveness in various attack scenarios. The results also demonstrate that IBD-PSC achieves a substantial improvement in detection performance compared to the defense baselines. We also provide the ROC curves of defenses against four representative attacks in Appendix I. In contrast, all baseline defenses fail in some cases (marked in red), especially under attacks involving subtle alterations across multiple pixels (e.g., Blend, WaNet) or physical attacks. This failure is primarily caused by their implicit assumptions about backdoors, such as sample-agnostic triggers and robustness against image preprocessing. We also provide the results with PreActResNet18 (He et al., 2016b) and MobileNet (Krizhevsky et al., 2009) architectures in Appendix J. Additionally, for more experimental results under other attack baselines, please refer to our Appendix K.
We also calculated the inference time of all methods under identical and ideal conditions for evaluating efficiency. For example, we assume that defenders will load all required models and images simultaneously (with more memory requirements compared to the vanilla model inference). This comparison is fair and reasonable since different defenses differ greatly in their mechanisms and requirements. Detailed settings can be found in Appendix L. As shown in  Figure 5, the efficiency of our IBD-PSC is on par with or even better than all baseline defenses. Even compared to no defense, the extra time is negligible, although IBD-PSC may increase some storage/computational consumptions.
@@ 5.3 Ablation Study
Impact of Scaling Factor $\omega$. IBD-PSC generates scaled models by amplifying the learnable parameter values of the selected BN layers with a fixed scalar $\omega$. We hereby explore its effects to our method. Specifically, we vary $\omega$ from 1 to 2 and calculate the AUROC and F1 scores of IBD-PSC against three representative backdoor attacks (i.e., BadNets, WaNet, and BATT) on CIFAR-10. As shown in Figure 6, in the initial phase, increasing $\omega$ can significantly improve both the AUROC and F1 scores against different backdoor attacks. Furthermore, the AUROC and F1 scores converge to nearly one and stabilize at approximately one for $\omega$ values of 1.5 or higher, i.e., the scaling factor has a relatively minor influence when it is sufficiently large.
Impact on Benign Samples within the Target Class. We evaluated the False Positive Rate (FPR) to assess the impact of our defense on benign samples within the target class. The results, shown in Table 4, demonstrate that although our defense, labeled ’Ours-C’, does incur some false positives, it significantly better than that of the variant with label consistency (dubbed ’Ours-L’) and SCALE-UP. This is primarily because the amplification process might also reduce the prediction confidence of benign samples from the target class, potentially due to ’short-cut’ classes that are easier to predict. We plan to explore ways to further mitigate this issue in our future work. Therefore, we used consistency of confidence rather than that of the predicted label (that used in SCALE-UP) for detection.
Table 4: The False Positive Rate (FPR) (%) of SCALE-UP and our defense on target and benign classes on CIFAR-10. In our cases, the target label is 0.
[…]
BadNets 72.74 29.00 87.73 10.73 0.20 1.88
Blend 54.28 19.80 22.55 3.39 18.34 2.64
PhysicalBA 90.58 23.98 60.72 19.20 4.10 1.50
WaNet 76.70 28.11 81.41 10.05 69.20 8.16
ISSBA 93.93 20.70 20.94 3.00 1.86 2.39
BATT 57.74 18.78 81.54 9.61 7.19 8.15
SRA 65.55 29.33 0.62 10.48 0.50 10.13
Additionally, we conduct comprehensive ablation studies as detailed in Appendix M to evaluate the robustness of our defense. The results in Sections M.1, M.2 and M.3 indicate that our defense is not sensitive to the hyperparameter selection, consistently achieving stable and promising performance (greater than 0.9) with hyperparameters close to our default settings (i.e., $n=5$, $T=0.9$, $\epsilon=0.6$, and $\omega=1.5$). The results presented in Section M.4 show that our defense proves effective across various target label selections, with both AUROC and F1 scores consistently nearing 1. Moreover, as shown in Section M.5, our defense remains effective even with a minimal dataset of benign samples (as few as 25 samples). We also validate that our defense maintains minimal impact on the regular functionality of benign models in Section M.6. Finally, we highlight the necessity of parameter amplification in Section M.7 by demonstrating that reducing the parameters of the BN layer is insufficient for detecting poisoned samples.
@@ 5.4 Resistance to Potential Adaptive Attacks
We initially assess the performance of IBD-PSC against attacks with low poisoning rates. This is because a small poisoning rate $\rho$ can prevent models from over-fitting triggers, thus weakening the association between triggers and target labels, as demonstrated in previous studies (Guo et al., 2023b; Qi et al., 2023). Specifically, we conduct attacks (BadNets, WaNet, and BATT) on the CIFAR-10 dataset with $\rho$ ranging from 0.02 to 0.1, ensuring the attack success rates exceed 80%. The results in Figure 7 consistently demonstrate the effectiveness of IBD-PSC, with AUROC and F1 scores consistently above 0.98 and 0.95, respectively. We also emphasize that our defense consistently outperforms baseline defenses across a range of poisoning rates and remains highly effective against attacks with small poisoning rates on the SubImageNet-200 dataset. For detailed discussions, please refer to Section N.1.
We further evaluate the robustness of IBD-PSC against potential adaptive attacks in the worst-case scenario where adversaries possess complete knowledge of our defense. The knowledge enables attackers to tailor adaptive attacks specifically designed to counteract the effects of parameter amplification. We explore one form of adaptive attack that maintains accurate predictions of benign samples even when subjected to model parameter amplification. Additionally, we design another form of adaptive attack that reduces the confidence with which poisoned samples are classified into the target classes by parameters-amplified models, which is presented in Section N.2. Typically, a vanilla backdoored model functions normally with benign samples but yields adversary-specific predictions when exposed to poisoned samples. The loss function for training such backdoored models is defined as follows:
where $\mathcal{L}(\cdot)$ represents the cross entropy loss function.
We introduce an adaptive loss item specifically designed to ensure that the benign samples are correctly predicted when subjected to model parameter amplification. This loss $\mathcal{L}_{\textrm{ada}}$ is defined as:
Similar to previous experiments, we also employ the three representative backdoor attacks to develop adaptive attacks on the CIFAR-10 dataset. Further details can be found in Section N.2Table 5 demonstrates the sustained robustness of our IBD-PSC across all cases. The effectiveness primarily originates from our adaptive layer selection strategy, which dynamically identifies BN layers for amplification, regardless of whether it is a vanilla or an adaptive backdoored model. The layers selected during the inference stage typically differ from those used in the training phase, enabling the IBD-PSC to effectively detect poisoned samples.
@@
5.5 A Closer Look to the Effectiveness of our Method
To gain deeper insights, we delve into the mechanisms of both SCALE-UP and our IBD-PSC. We utilize t-SNE (Van der Maaten & Hinton, 2008) for visualizing the features of benign and poisoned samples in the last hidden layer. We adopt the representative BadNets attack method on the CIFAR-10 dataset as an example for our discussions. More results about other attack methods can be found in Appendix O. The results in Figure 8 demonstrate that both SCALE-UP and our IBD-PSC induce more significant shifts in the feature space for benign samples compared to the poisoned samples. These larger shifts result in changes in the predictions for benign samples. These results provide clear evidence of the effectiveness of the two defense methods. Furthermore, in contrast to SCALE-UP, our IBD-PSC method induces more significant shifts in benign samples. This disparity in shift magnitude may stem from the constrained pixel value range of [0, 255], potentially mitigating the impact of amplification. However, the values of model parameters do not have such bounded constraints. Consequently, the larger shifts contribute to a more distinct separation between benign and poisoned samples, significantly augmenting the effectiveness of IBD-PSC.
@@ 6 Conclusion
7 Statement of the Broader Impact
@@
Theorem 3.1. Let $F=FC\circ f_{L}\circ\dots\circ f_{1}$ be a backdoored DNN with $L$ hidden layers and FC denotes the fully-connected layers. Let $x$ be an input, $\bm{b}=f_{l}\circ\cdots\circ f_{1}(x)$ be its batch-normalized feature after the $l$-th layer ($1\leq l\leq L$), and $t$ represent the attacker-specified target class. Assume that $\bm{b}$ follows a mixture of Gaussian distributions. Then the following two statements hold: (1) Amplifying the ${\bm{\beta}}$ and $\gamma$ parameters of the $l$-th BN layer can make $\|\tilde{\bm{b}}\|_{2}$ ($\tilde{\bm{b}}$ is the amplified version of $\bm{b}$) arbitrarily large, and (2) There exists a positive constant $M$ that is independent of $\tilde{\bm{b}}$, such that whenever $\|\tilde{\bm{b}}\|_{2}>M$, then $\arg\max FC\circ f_{L}\circ\dots\circ f_{l+1}(\tilde{\bm{b}})=t$, even when $\arg\max FC\circ f_{L}\circ\dots\circ f_{l+1}(\bm{b})\not=t$
@@
We assume all features follow the mixture of Gaussians. The assumption is commonly used in many deep learning theory papers (Guo et al., 2024; Zoran & Weiss, 2012; Loureiro et al., 2021) as it simplifies the analysis and provides a tractable framework for modeling complex data distributions. So ${\bm{a}}$ and $\tilde{{\bm{a}}}$ will follow:
@@
For a sufficiently trained network, it is well-known that, with the neural collapse (Papyan et al., 2020), $\mu_{c}$ and $\sigma_{c},c\in\mathcal{Y}$ form a simplex and are uniformly distributed. Below we will try to find out the characteristics of a backdoored model.
@@ A.1 Characterize the Backdoored Model
We assume that the trigger $\bm{t}\in\mathcal{X}$, which fools the backdoored model $\tilde{\mathcal{F}}$ to recognize ${\bm{x}}+\bm{t}$ as the attacked target class $t$ instead of its true class $c$, is very small, i.e., $\|\bm{t}\|_{2}\ll 1$. In this paper, we assume that all images have been normalized, i.e., $x\in[0,1]$. Accordingly, $||t||_{2}\ll 1$ holds true in practice since the triggers are either very sparse (e.g., BadNets) or have a small overall magnitude (e.g., WaNet). So the feature distribution of ${\bm{x}}(\bm{t})={\bm{x}}+\bm{t}$ may be approximated by
@@
As ${\bm{x}}(\bm{t})$ should be recognized as category $t$, the conditional probability of $\tilde{\bm{b}}(\bm{t})$ being sampled from $\tilde{\bm{b}}|c$ should be smaller than from $\tilde{\bm{b}}|t$ for all $c\in\mathcal{Y}$, and we thus can get, $\forall{\bm{x}},\bm{t}\in\mathcal{X}$,
@@ Appendix B Detailed Configurations of the Empirical Study in Section 3
In this section, we adopt BadNets (Gu et al., 2017), WaNet (Nguyen & Tran, 2021), and BATT (Xu et al., 2023) as examples for our analysis. These attacks epitomize static, dynamic, and physical backdoor attacks, respectively. Our experiments are conducted on the CIFAR-10 dataset (Krizhevsky et al., 2009), using the ResNet18 model (He et al., 2016a). For each attack, we set the poisoning rate ($\rho$) to 0.1, achieving attack success rates over 99%. In particular, we implement the backdoor attacks using their official codes with default settings. Specifically, the backdoor trigger for BadNets is represented as a $3\times 3$ grid in black-and-white and is added to the lower-right corner of the poisoned images. For WaNet, the trigger is applied to the original images through elastic image warping transformation. In the case of BATT, the poisoned samples are obtained by rotating the original images by sixteen degrees. These attacks are implemented using the BackdoorBox toolkit (Li et al., 2023) [^1] https://github.com/THUYimingLi/BackdoorBox.
Regarding the scaling procedure, we adopt a layer-wise weight scaling operation to generate the parameter-amplified models. we scale up on the BN parameters (i.e., $\gamma$ and ${\bm{\beta}}$) with $\omega=1.5$ times starting from the last layer and gradually moving forward to more layers. For example, in a 20-layer model, the first iteration involves scaling the weights of the 20th layer, and the next iteration extends the scaling to the 20th and the 19th layers, and so on. We then calculate the average confidence of 2000 testing samples for each parameter-scaled model. The average confidence is defined as the average probability of samples on the label predicted by the original unamplified model.

### v3

@@ 5.2 Main Results
As shown in Table 1-3, IBD-PSC consistently achieves promising performance in all cases across various datasets. For instance, it achieves AUROC and F1 scores approaching 1.0, indicating its effectiveness in various attack scenarios. The results also demonstrate that IBD-PSC achieves a substantial improvement in detection performance compared to the defense baselines. In contrast, all baseline defenses fail in some cases (marked in red), especially under attacks involving subtle alterations across multiple pixels (e.g., Blend, WaNet) or physical attacks. This failure is primarily caused by their implicit assumptions about backdoors, such as sample-agnostic triggers and robustness against image preprocessing. We also provide the results with PreActResNet18 (He et al., 2016b) and MobileNet (Krizhevsky et al., 2009) architectures in Appendix I. We also provide the ROC curves of defenses against four representative attacks in Appendix J. Besides, for more experimental results under other attack baselines in Appendix K.
We also calculated the inference time of all methods under identical and ideal conditions for evaluating efficiency. For example, we assume that defenders will load all required models and images simultaneously (with more memory requirements compared to the vanilla model inference). Arguably, this comparison is fair and reasonable since different defenses differ greatly in their mechanisms and requirements. Detailed settings can be found in Appendix L. As shown in  Figure 5, the efficiency of our IBD-PSC is on par with or even better than all baseline defenses. The extra time is negligible compared to no defense, although IBD-PSC may increase some storage or computational consumptions. More detailed discussions about our running efficiency and storage requirements are in Appendix R.
@@ 5.3 Ablation Study
Impact of Scaling Factor $\omega$. IBD-PSC generates scaled models by amplifying the learnable parameter values of the selected BN layers with a fixed scalar $\omega$. We hereby explore its effects on our method. Specifically, we vary $\omega$ from 1 to 2 and calculate the AUROC and F1 scores of IBD-PSC against three representative backdoor attacks (i.e., BadNets, WaNet, and BATT) on CIFAR-10. As shown in Figure 6, in the initial phase, increasing $\omega$ can significantly improve both the AUROC and F1 scores against different backdoor attacks. Furthermore, the AUROC and F1 scores converge to nearly one and stabilize at approximately one for $\omega$ values of 1.5 or higher, i.e., the scaling factor has a relatively minor influence when it is sufficiently large. Besides, we conduct further ablation studies on other hyper-parameters of our method, as detailed in Appendix M.
Impact of Confidence Consistency. IBD-PSC leverages the consistency of confidence for detection, in contrast to the SCALE-UP method, which relies on the consistency of the predicted label. SCALE-UP is designed for black-box scenarios where defenders only have access to predicted labels, while our IBD-PSC focuses on white-box settings where predicted confidences are naturally available. To validate the effectiveness of IBD-PSC, we develop a variant that uses label consistency (dubbed ‘Ours-L’). We then calculate the False Positive Rate (FPR) (%) for both target and benign classes on the CIFAR-10 dataset across various backdoor attacks. As shown in Table 4, our method significantly reduces false positives in both the target and benign classes, outperforming both the Ours-L and SCALE-UP.
@@ 5.4 Resistance to Potential Adaptive Attacks
We initially assess the performance of IBD-PSC against attacks with low poisoning rates. This is because a small poisoning rate $\rho$ can prevent models from over-fitting triggers, thus weakening the association between triggers and target labels, as demonstrated in previous studies (Guo et al., 2023b; Qi et al., 2023). Specifically, we conduct attacks (BadNets, WaNet, and BATT) on the CIFAR-10 dataset with $\rho$ ranging from 0.02 to 0.1, ensuring the attack success rates exceed 80%. The results in Figure 7 consistently demonstrate the effectiveness of IBD-PSC, with AUROC and F1 scores consistently above 0.98 and 0.95, respectively. Results on SubImageNet-200 are shown in Section N.1.
Table 4: The False Positive Rate (FPR) (%) of SCALE-UP and our defense on target and benign classes on the CIFAR-10 dataset.
[…]
BadNets 72.74 29.00 0.40 9.76 0.20 1.88
Blend 54.28 19.80 22.55 3.39 18.34 2.64
PhysicalBA 90.58 23.98 4.60 5.42 4.10 1.50
WaNet 76.70 28.11 81.41 10.05 69.20 8.16
ISSBA 93.93 20.70 20.94 3.00 17.22 0.61
BATT 57.74 18.78 2.35 9.72 0.87 6.90
SRA 65.55 29.33 0.62 10.48 0.50 10.13
Ada-Patch 93.80 25.77 8.67 4.78 4.34 3.00
@@
(a) BadNets
(b) Blend
(b) BATT
(f) Ada-patch
Figure 8: The violin plots of the prediction confidences for benign samples in the target and other classes, as well as for poisoned samples, as predicted by the initial and scaled models on CIFAR-10. The threshold is 0.9.
We further evaluate the robustness of IBD-PSC against potential adaptive attacks in the worst-case scenario, where adversaries possess complete knowledge of our defense. Typically, a vanilla backdoored model functions normally with benign samples but yields adversary-specific predictions when exposed to poisoned samples. The loss function for training such backdoored models is defined as follows:
where $\mathcal{L}(\cdot)$ represents the cross entropy loss function.
We design an adaptive loss term $\mathcal{L}_{\textrm{ada}}$ to ensure benign samples are correctly predicted under parameter amplification:
Similar to previous experiments, we also employ the three representative backdoor attacks to develop adaptive attacks on the CIFAR-10 dataset. Table 5 demonstrates the sustained robustness of our IBD-PSC across all cases. The effectiveness primarily originates from our adaptive layer selection strategy, which dynamically identifies BN layers for amplification, regardless of whether it is a vanilla or an adaptive backdoored model. The layers selected during the inference stage typically differ from those used in the training phase, enabling the IBD-PSC to effectively detect poisoned samples. More results and the resistance to another adaptive attack can be found in Section N.2.
5.5 Performance on Benign Samples from Target Class
In this section, we evaluate the effectiveness of our defense on benign samples from the target class. We conduct experiments on the CIFAR-10 dataset against four different attacks and present the confidences of both the initial backdoored model and the scaled models in Figure 8. As shown, the confidences of benign samples from both the target class and other classes decrease due to parameter amplification, falling below the threshold. In contrast, the confidence values for poisoned samples mostly remain above this threshold. These results demonstrate that our defense effectively distinguishes between benign and poisoned samples, regardless of whether the benign sample originates from the target class. In particular, we observe an interesting phenomenon that scaled models tend to cluster the confidences for benign samples from the target class in the more difficult-to-learn class(es), rather than in the easier ones, which is unexpected. We will further explore its intrinsic mechanism in our future work. Additional analysis can be found in Appendix O.
5.6 A Closer Look to the Effectiveness of our Method
To gain deeper insights, we delve into the mechanisms of both SCALE-UP and our IBD-PSC. We utilize t-SNE (Van der Maaten & Hinton, 2008) for visualizing the features of benign and poisoned samples in the last hidden layer. We adopt the representative BadNets attack method on the CIFAR-10 dataset as an example for our discussions. More results about other attack methods can be found in Appendix P. The results in Figure 9 demonstrate that both SCALE-UP and our IBD-PSC induce more significant shifts in the feature space for benign samples compared to the poisoned samples. These larger shifts result in changes in the predictions for benign samples. These results provide clear evidence of the effectiveness of the two defense methods. Furthermore, in contrast to SCALE-UP, our IBD-PSC method induces more significant shifts in benign samples. This disparity in shift magnitude may stem from the constrained pixel value range of [0, 255], potentially mitigating the impact of amplification. However, the values of model parameters do not have such bounded constraints. Consequently, the larger shifts contribute to a more distinct separation between benign and poisoned samples, significantly augmenting the effectiveness of IBD-PSC.
5.7 The Extension to Training Set Purification
Although our method is initially and primarily designed to filter malicious testing samples, it can also be used to detect potentially poisoned samples within a compromised training set. Specifically, users can first train a model on this dataset with a standard process and then exploit our detection method. To verify our effectiveness, we conduct experiments on the CIFAR-10 dataset against three representative attacks. The results show a 100% TPR and nearly 100% AUROC scores, with FPR scores close to 0%. We compare the detection performance of our method with the most advanced defenses, i.e., CD (Huang et al., 2023) and MSPC (Pal et al., 2024), and the results show that our method achieves the best detection performance. Subsequently, we retrain a model on this purified dataset to evaluate both its BA and the ASR. The ASR scores of these retrained models are less than 0.5%, rendering the attacks ineffective. More settings and results can be found in Appendix Q.
@@ 6 Conclusion
Acknowledgement
This work was supported in part by the National Natural Science Foundation of China under Grants 62071142 and by the Guangdong Basic and Applied Basic Research Foundation under Grant 2024A1515012299.
Impact Statement
@@ Acknowledgement
Theorem 3.1. Let $F=FC\circ f_{L}\circ\dots\circ f_{1}$ be a backdoored DNN with $L$ hidden layers and FC denotes the fully-connected layers. Let $x$ be an input, $\bm{b}=f_{l}\circ\cdots\circ f_{1}(x)$ be its batch-normalized feature after the $l$-th layer ($1\leq l\leq L$), and $t$ represent the attacker-specified target class. Assume that $\bm{b}$ follows a mixture of Gaussian distribution. Then the following two statements hold: (1) Amplifying the ${\bm{\beta}}$ and $\gamma$ parameters of the $l$-th BN layer can make $\|\tilde{\bm{b}}\|_{2}$ ($\tilde{\bm{b}}$ is the amplified version of $\bm{b}$) arbitrarily large, and (2) There exists a positive constant $M$ that is independent of $\tilde{\bm{b}}$, such that whenever $\|\tilde{\bm{b}}\|_{2}>M$, then $\arg\max FC\circ f_{L}\circ\dots\circ f_{l+1}(\tilde{\bm{b}})=t$, even when $\arg\max FC\circ f_{L}\circ\dots\circ f_{l+1}(\bm{b})\not=t$
@@
We assume all features follow the mixture of Gaussians, an assumption commonly used in many deep learning theory papers (Guo et al., 2024; Zoran & Weiss, 2012; Loureiro et al., 2021) as it simplifies analysis and provides a tractable framework for modeling complex data distributions. Consequently, ${\bm{a}}$ and $\tilde{{\bm{a}}}$ follow:
@@
For a sufficiently trained network, it is well-known that, with the neural collapse (Papyan et al., 2020), $\mu_{c}$ and $\sigma_{c},c\in\mathcal{Y}$ form a simplex and are uniformly distributed. Specifically, in neural collapse scenarios, the features of each class form a simplex equiangular tight frame. This means that all features share (nearly) the same within-class variance and exhibit uniform mean values. Below, we try to find out the characteristics of a backdoored model.
@@ A.1 Characterize the Backdoored Model
We denote the poisoned sample as $x(t)=x+G(x)$, where $x$ is a benign input and $G(x)$, the trigger, is very small (i.e., $\|G(x)\|_{2}\ll 1$). The trigger $G(x)$ can be static or vary with different inputs, which fools the backdoored model $\tilde{\mathcal{F}}$ into recognizing the poisoned samples as the attacked target class $t$ instead of its true class $c$. For clarity, we simplify the trigger $G(x)$ as $\bm{t}$. In this paper, we assume that all images have been normalized, i.e., $x\in[0,1]$. Accordingly, $||t||_{2}\ll 1$ holds in practice since the triggers are either very sparse (e.g., BadNets) or have a small overall magnitude (e.g., WaNet). So the feature distribution of ${\bm{x}}(\bm{t})={\bm{x}}+\bm{t}$ may be approximated by
@@
As ${\bm{x}}(\bm{t})$ should be recognized as category $t$, the conditional probability of $\tilde{\bm{b}}(\bm{t})$ being sampled from $\tilde{\bm{b}}|c$ should be smaller than from $\tilde{\bm{b}}|t$ for all $c\in\mathcal{Y}$. The assumption holds, particularly for the deeper hidden layers, under the Gaussian mixture distribution and a well-trained network. Specifically, in Equations A3 and A4, we assume the conditional distribution $a_{l}(x)|\arg\max f(x)=c$ to be Gaussian. This conditional distribution is derived after completing the forward pass and examining the previous layers, and it remains unchanged, i.e., $p(a_{l}(x)\in a_{l}(A)|\arg\max f(x)=c)$$=p(a_{l+1}(x)\in a_{l+1}(A)|\arg\max f(x)=c),A={x:\arg\max f(x)=s}$. Clearly, in the last layer, the probability of $a_{l}(x)$ belonging to class $c$ will always be larger than other classes. Therefore, the assumption holds for the conditional distribution of $a_{l}(x)|\arg\max f(x)=c$. Thus, we can get, $\forall{\bm{x}},\bm{t}\in\mathcal{X}$,
@@ Appendix B Detailed Configurations of the Empirical Study in Section 3
In this section, we adopt BadNets (Gu et al., 2017), WaNet (Nguyen & Tran, 2021), and BATT (Xu et al., 2023) as examples for our analysis. These attacks epitomize static, dynamic, and physical backdoor attacks, respectively. Our experiments are conducted on the CIFAR-10 dataset (Krizhevsky et al., 2009), using the ResNet18 model (He et al., 2016a). For each attack, we set the poisoning rate ($\rho$) to 0.1, achieving ASRs over 99%. In particular, we implement the backdoor attacks using their official codes with default settings. Specifically, the backdoor trigger for BadNets is represented as a $3\times 3$ grid in black-and-white and is added to the lower-right corner of the poisoned images. For WaNet, the trigger is applied to the original images through elastic image warping transformation. In the case of BATT, the poisoned samples are obtained by rotating the original images by sixteen degrees. These attacks are implemented using the BackdoorBox toolkit (Li et al., 2023a) [^1] https://github.com/THUYimingLi/BackdoorBox.
Regarding the scaling procedure, we adopt a layer-wise weight scaling operation to generate the parameter-amplified models. we scale up on the BN parameters (i.e., $\gamma$ and ${\bm{\beta}}$) with $\omega=1.5$ times starting from the last layer and gradually moving forward to more layers. For example, in a 20-layer model, the first iteration involves scaling the weights of the 20th layer, and the next iteration extends the scaling to the 20th and the 19th layers, and so on. We then calculate the average confidence of 2000 testing samples for each parameter-scaled model. In this paper, confidence refers to the predicted probability assigned to an input sample for a specified label. For instance, if an image of a cat is predicted as the cat label with a probability of 0.9, then the confidence of the input under the cat label is 0.9. The average confidence is defined as the average probability of samples on the label predicted by the original unamplified model.